# A Bias–Variance Decomposition for Ensembles over Multiple Synthetic Datasets

### v2

@@ 3 Experiments
In this section, we describe our experiments. The common theme in all of them is generating synthetic data and evaluating the performance of several downstream prediction algorithms trained on the synthetic data. The performance evaluation uses a test set of real data, which is split from the whole dataset before generating synthetic data.
@@ 3.1 Effect of Multiple Synthetic Datasets
Figure 1: MSE on regression datasets (a) or Brier score on classification datasets (b) of the ensemble of downstream predictors, with varying number of synthetic datasets $m$ from synthpop. Increasing the number of synthetic datasets generally decreases both metrics, especially for decision trees and 1-NN. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM), ridge regression (RR) and logistic regression (LogR). The black line is the MSE of the best predictor on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Tables S4 to S10 in the Appendix contain the numbers from the plots.
As our first experiment, we evaluate the performance of the synthetic data ensemble on 7 datasets, containing 4 regression and 3 classification tasks. See Appendix D.1 for details on the datasets. We use the synthetic data generators DDPM (Kotelnikov et al., 2023) and synthpop (Nowok et al., 2016), which we selected after a preliminary experiment described in Appendix E.1. We only plot the results from synthpop in the main text to save space, and defer the results of DDPM to Figure S2 in the Appendix. We generate 32 synthetic datasets, of which between 1 and 32 are used to train the ensemble. The results are averaged over 3 runs with different train-test splits.
@@
In Appendix E.4, we estimate the MV and SDV terms of the decomposition in Theorem 2.2. These confirm our claims on the model variances: decision trees and 1-NN have a high variance, while linear, ridge and logistic regression have a low variance.
(a) ACS 2018
(b) German credit
Figure 2: MSE or Brier score prediction on two datasets. The predictions are very accurate on these datasets. The solid lines for DDPM and synthpop (SP-P) show the same error MSE or Brier score as Figure 1, while the dashed lines show predicted MSE or Brier score. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. We omitted downstream algorithms with uninteresting flat curves from panel (a). See Figure S6 in the Appendix for the full panel (a), and Figures S7 and S8 for the other datasets. Tables S11 and S12 in the Appendix contain the numbers from the plots. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@ 3.2 Predicting performance from Two Synthetic Datasets
Figure 2 contains the results for the ACS 2018 and German credit datasets, and Figures S6 to S8 in the Appendix contain the results for the other datasets. The predictions are very accurate on ACS 2018 and German credit, and reasonably accurate on the other datasets. The variance of the prediction depends heavily on the variance of the errors computed from the test data.
@@ 3.3 Differentially Private Synthetic Data
Figure 3: Brier score of the ensemble of downstream predictors with varying numbers of synthetic datasets $m$, generated with the DP methods AIM or NAPSU-MQ from the Adult dataset with a reduced set of features. Increasing the number of datasets generally decreases the score, even with AIM which splits the privacy budget between $m$ synthetic datasets. The privacy parameters are $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$. The predictors are the same as in Figure 1. The black lines show the loss of the best non-DP downstream predictor trained on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Table S13 contains the numbers from the plots, and Figure S10 contains plots of the other error metrics.
@@
The results for Brier score are in Figure 3, and the results for all metrics are in Figure S10 in the Appendix. Increasing $m$ almost always improves the results, which is expected for NAPSU-MQ due to Theorem 2.4, but somewhat surprising for AIM, which splits the privacy budget. Once again, the high-variance predictors 1-NN, 5-NN and decision trees benefit the most from multiple synthetic datasets. AIM always outperforms NAPSU-MQ, which is likely a result of AIM being able to handle a more complicated set of input queries than NAPSU-MQ.
@@ 4 Discussion
Limitations
Our main theory assumes that the synthetic datasets are generated i.i.d. given either the real data or a DP summary. We generalise the bias-variance decompositions to non-i.i.d. settings in Appendix B. The implications of the decompositions, like the rule-of-thumb on the number of synthetic datasets, may not apply if i.i.d. assumption is removed.
Broader Impacts
Generating multiple synthetic datasets can potentially increase disclosure risk if appropriate precautions, such as differential privacy, are not used. Because of this, care should be taken and precautions like differential privacy should be used if possible when releasing multiple synthetic datasets of sensitive data, despite our theory suggesting that multiple synthetic datasets are beneficial for utility.
Conclusion
We derived bias-variance decompositions for using synthetic data in several cases: for MSE or Brier score with i.i.d. synthetic datasets given the real data (Section 2.2) and MSE with i.i.d. synthetic datasets given a DP summary of the real data (Section 2.4). We generalised these decompositions to non-i.i.d. synthetic datasets (Appendix B) and Bregman divergences (Appendix C). The decompositions make actionable predictions, such as yielding a simple rule of thumb that can be used to select the number of synthetic datasets. We empirically examined the performance of generative ensembles on several real datasets and downstream predictors, and found that the predictions of the theory generally hold in practice (Section 3). These findings significantly increase the theoretical understanding of generative ensembles, which is very limited in prior literature.
Acknowledgements
@@ Appendix A Missing Proofs
$\mathrm{MV}$ $=\operatorname*{\mathbb{E}}_{D_{r},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]$ (42)
$\mathrm{SDV}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]$ (43)
$\mathrm{COV}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\left[\operatorname*{Cov}_{D_{s}^{1},D_{s}^{2}|D_{r}}[g(x;D_{s}^{1}),g(x;D_{s}^{2})]\right]$ (44)
$\mathrm{RDV}$ $=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]$ (45)
$\mathrm{SDB}$ $=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|D_{r}}[f_{\theta}(x)]]$ (46)
$\mathrm{MB}$ $=\operatorname*{\mathbb{E}}_{D_{r},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].$ (47)
@@ C.1 Background: Bregman Divergences
where $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is a strictly convex differentiable function. Many common error metrics, like MSE and cross entropy, can be expressed as expected values of a Bregman divergence. In fact, proper scoring rules555 Proper scoring rules are error metrics that are minimised by predicting the correct probabilities. can be characterised via Bregman divergences (Gneiting & Raftery, 2007; Kimpara et al., 2023). Table S1 shows how the metrics we consider are expressed as Bregman divergences (Gupta et al., 2022).
@@
Brier Score (2 classes) $(y_{1}-p_{1})^{2}$ $t^{2}$ $\frac{1}{m}\sum_{j=1}^{m}p^{(j)}$
@@ C.2 Bregman Divergence Decomposition for Synthetic Data
This decomposition is not as easy to interpret as the other two in Section 2, as is only gives an upper bound, and does not explicitly depend on the number of synthetic datasets.
@@ D.1 Datasets
In our experiments, we use 7 tabular datasets. For four of them, the downstream prediction task is regression, and for the other three, the prediction task is binary classification. Table S2 lists some general information on the datasets. We use 25% of the real data as a test set, with the remaining 75% being used to generate the synthetic data, for all of the datasets. All experiments are repeated several times, with different train-test splits for each repeat.
@@
(Nash et al., 1995, CC BY 4.0) The abalone dataset contains information on abalones, with the task of predicting the number of rings on the abalone from the other information.
@@
(Kohavi & Becker, 1996, CC BY 4.0) This dataset is the UCI Adult dataset with a reduced set of features. The subset of features is age, workclass, education, marital-status, race, gender, capital-gain, capital-loss and hours-per-week. [^6] This subset was used in the NAPSU-MQ experiments of Räisä et al. (2023b) We binarise capital-gain and capital-loss to indicate whether the original value is positive or not for both synthetic data generation and downstream prediction. We discretise age and hours-per-week to 5 categories for synthetic data generation, and converted back to continuous values for the downstream prediction, as our synthetic data generators require discrete data. We only remove rows with missing values in the included columns, so the number of rows is larger with the reduced set of features.
@@
(https://scikit-learn.org/stable/datasets/real_world.html#california-housing-dataset, license not mentioned) The california housing dataset contains information on housing districts, specifically census block groups, in California. The task is predicting the median house value in the district. We removed outlier rows where the average number of rooms is at least 50, or the average occupancy is at least 30. According to the dataset description, these likely correspond to districts with many empty houses. We log-transformed the target variable, as well as the population and median income features.
@@ E.2 Extra Plots
Figure S6: MSE prediction on the first two regression datasets. The linear regression measured MSE line for synthpop with ACS 2018 data is cut off due to excluding repeats with extremely large MSE ($\geq 10^{6}$). 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure S7: MSE prediction on the last two regression datasets. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure S8: Brier score prediction on two classification datasets. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.

### v3

@@ 3 EXPERIMENTS
In this section, we describe our experiments. The common theme in all of them is generating synthetic data and evaluating the performance of several downstream prediction algorithms trained on the synthetic data. The performance evaluation uses a test set of real data, which is split from the whole dataset before generating synthetic data. Our code is available at https://github.com/DPBayes/generative-ensemble-bias-variance-decomposition.
@@
In each of our figures comparing prediction performance, we have included a horizontal black line showing the performance of the best downstream predictor without synthetic data. The downstream models include a random forest and gradient boosted trees, which are ensemble methods, so this line serves as a baseline for ensembles without synthetic data. Another relevant baseline is the performance of an ensemble with only one synthetic dataset, which is given by the results of random forests and gradient boosted trees with $m=1$.
@@ 3.1 Effect of Multiple Synthetic Datasets
Figure 1: MSE on regression datasets (a) or Brier score on classification datasets (b) of the ensemble of downstream predictors, with varying number of synthetic datasets $m$ from synthpop. Increasing the number of synthetic datasets generally decreases both metrics, especially for decision trees and 1-NN. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM), ridge regression (RR) and logistic regression (LogR). The black line is the MSE of the best predictor on real data. Tables S4 to S10 in the Appendix contain the numbers from the plots.
As our first experiment, we evaluate the performance of the synthetic data ensemble on 7 datasets, containing 4 regression and 3 classification tasks. See Appendix D.1 for details on the datasets. We use the synthetic data generators DDPM (Kotelnikov et al., 2023) and synthpop (Nowok et al., 2016), which we selected after a preliminary experiment described in Appendix E.1. We only plot the results from synthpop in the main text to save space, and defer the results of DDPM to Appendix E.2. We generate 32 synthetic datasets, of which between 1 and 32 are used to train the ensemble. The results are averaged over 3 runs with different train-test splits. We compute error bars as 95% confidence intervals obtained from bootstrapping over the repeats.
@@
In Figure 2, we estimate the MV and SDV terms of the decomposition in Theorem 2.2. The results show that MV depends mostly on the downstream predictor, while SDV also depends on the synthetic data generator. The results also confirm our claims on the model variances: decision trees and 1-NN have a high variance, while linear, ridge and logistic regression have a low variance. See Appendix E.4, for details, and Figure S11 for results on all datasets.
(a) Abalone
(b) Adult
Figure 2: Estimating the MV and SDV terms from the decomposition. Decision trees have high variances on all datasets, while linear, ridge and logistic regression have low variances. MV depends mostly on the predictor, while SDV depends on both the predictor and synthetic data generation algorithm. The points are the averages of estimated MV and SDV, averaged over the test data, from 3 repeats with different train-test splits. See Figure S11 in the Appendix for results on all datasets.
In Appendix E.5 we compare an alternative to the ensemble of multiple synthetic datasets: generating a single large synthetic datasets with an equal number of datapoints as all the multiple synthetic datasets combined. One could expect generating a larger synthetic dataset to also reduce variances while saving on the computational cost of training multiple generative models. However, in the cases we examined in Figure S12, a single larger dataset gave at best a small improvement, and sometimes even increased the error. In contrast, adding more synthetic datasets often decreases error and never increases it. As a result, as a default choice, we recommend generating as many synthetic datasets as possible that have the same size as the original. We do not recommend making the synthetic datasets smaller, since the computational saving is likely small, and the additional generators that could be trained with the savings have diminishing returns due to the rule-of-thumb that $m$ synthetic datasets give a $(1-\frac{1}{m})$ fraction of the benefit.
Figure 3: MSE or Brier score prediction on the ACS 2018 dataset. The predictions are very accurate on this dataset. The solid lines for DDPM and synthpop (SP-P) show the same error MSE or Brier score as Figure 1, while the dashed lines show predicted MSE or Brier score. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. We omitted downstream algorithms with uninteresting flat curves. See Figure S6 in the Appendix for the full figure, and Figures S7 and S8 for the other datasets. Tables S11 and S12 in the Appendix contain the numbers from the plots.
@@ 3.2 Predicting Performance from Two Synthetic Datasets
Figure 3 contains the results for the ACS 2018 datasets, and Figures S6 to S8 in the Appendix contain the results for the other datasets. The predictions are very accurate on ACS 2018, and reasonably accurate on the other datasets. The variance of the prediction depends heavily on the variance of the errors computed from the test data.
@@ 3.3 Differentially Private Synthetic Data
Figure 4: Brier score of the ensemble of downstream predictors with varying numbers of synthetic datasets $m$, generated with the DP methods AIM or NAPSU-MQ from the Adult dataset with a reduced set of features. Increasing the number of datasets generally decreases the score, even with AIM, which splits the privacy budget between $m$ synthetic datasets. The privacy parameters are $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$. The predictors are the same as in Figure 1. The black lines show the loss of the best non-DP downstream predictor trained on real data. Table S13 contains the numbers from the plots, and Figure S10 contains plots of the other error metrics.
@@
The results for Brier score are in Figure 4, and the results for all metrics are in Figure S10 in the Appendix. Increasing $m$ often always improves the results, which is expected for NAPSU-MQ due to Theorem 2.4, but somewhat surprising for AIM, which splits the privacy budget. For AIM, $m=4$ looks like a sensible default choice for the well-performing downstream predictors, since performance does not greatly change either way with $m>4$. Once again, the high-variance predictors 1-NN, 5-NN and decision trees benefit the most from multiple synthetic datasets. AIM always outperforms NAPSU-MQ, which is likely a result of AIM being able to handle a more complicated set of input queries than NAPSU-MQ.
@@ 4 DISCUSSION
Limitations. Our main theory assumes that the synthetic datasets are generated i.i.d. given either the real data or a DP summary, which for example leaves out generative ensembles that explicitly encourage diversity between synthetic datasets in some way. We generalise the bias-variance decompositions to non-i.i.d. settings in Appendix B. The implications of the decompositions, like the rule-of-thumb on the number of synthetic datasets, do not apply in general if i.i.d. assumption is removed, but they can still apply with additional assumptions. We give an example in Appendix B.
Conclusion. We derived bias-variance decompositions for using synthetic data in several cases: for MSE or Brier score with i.i.d. synthetic datasets given the real data (Section 2.2) and MSE with i.i.d. synthetic datasets given a DP summary of the real data (Section 2.4). We generalised these decompositions to non-i.i.d. synthetic datasets (Appendix B) and Bregman divergences (Appendix C). The decompositions make actionable predictions, such as yielding a simple rule of thumb that can be used to select the number of synthetic datasets. We empirically examined the performance of generative ensembles on several real datasets and downstream predictors, and found that the predictions of the theory generally hold in practice (Section 3). These findings significantly increase the theoretical understanding of generative ensembles, which is very limited in prior literature.
Acknowledgments
@@
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Not Applicable]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [No] None of the providers of the datasets we used provide information about consent.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Appendix A MISSING PROOFS
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{COV}&=\operatorname*{\mathbb{E}}_{D_{r}}\left[\operatorname*{Cov}_{D_{s}^{1},D_{s}^{2}|D_{r}}[g(x;D_{s}^{1}),g(x;D_{s}^{2})]\right]\\ \mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|D_{r}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].\end{split}$ (43)
@@
Implication of Non-i.i.d.  Theory
From Theorems B and B, it is clear that the implications of the i.i.d. theory do not always hold. For example, when increasing the number of synthetic datasets with each generator being a different model, all of the terms in Theorems B and B that are averages over the generators can change.
However, we can recover some implications with additional assumptions. For example, if we assume that the generators and downstream predictors are always the same, possibly correlated, but the covariance term $\mathrm{COV}$ does depend on the number of synthetic datasets, we can derive a similar MSE prediction rule as in Section 2.3 from Equation (42).
@@ C.1 Background: Bregman Divergences
where $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is a strictly convex differentiable function. Many common error metrics, like MSE and cross entropy, can be expressed as expected values of a Bregman divergence. In fact, proper scoring rules666 Proper scoring rules are error metrics that are minimised by predicting the correct probabilities. can be characterised via Bregman divergences (Gneiting & Raftery, 2007; Kimpara et al., 2023). Table S1 shows how the metrics we consider are expressed as Bregman divergences (Gupta et al., 2022).
@@
Brier Score (2 classes) $(y_{0}-p_{0})^{2}$ $t^{2}$ $\frac{1}{m}\sum_{j=1}^{m}p^{(j)}$
@@ C.2 Bregman Divergence Decomposition for Synthetic Data
This decomposition is not as informative as the other two in Section 2, as is only gives an upper bound, and does not explicitly depend on the number of synthetic datasets.
Implication for Non-Synthetic Data Ensembles
The theory of Gupta et al. (2022) assumes that each member of the ensemble is trained with an independently sampled real dataset. In practice, this would mean that one needs to split the training data between each ensemble member to apply their theory, so their theory does not apply to any of the ways ensembles are usually trained. In contrast, our theory applies to bagging, as discussed in Section 2.1, so our Theorem C.2 implies a generalisation of Proposition 5.3 of Gupta et al. (2022) to bagging.
@@ D.1 Datasets
In our experiments, we use 7 tabular datasets. For four of them, the downstream prediction task is regression, and for the other three, the prediction task is binary classification. Table S2 lists some general information on the datasets. We use 25% of the real data as a test set, with the remaining 75% being used to generate the synthetic data, for all of the datasets. All experiments are repeated several times, with different randomly selected train-test splits for each repeat.
@@
(Nash et al., 1995, CC BY 4.0) The abalone dataset contains information on abalones, with the task of predicting the number of rings on the abalone from other information like weight and size.
@@
(Kohavi & Becker, 1996, CC BY 4.0) This dataset is the UCI Adult dataset with a reduced set of features. The subset of features is age, workclass, education, marital-status, race, gender, capital-gain, capital-loss and hours-per-week. [^7] This subset was used in the NAPSU-MQ experiments of Räisä et al. (2023b) We binarise capital-gain and capital-loss to indicate whether the original value is positive or not for both synthetic data generation and downstream prediction. We discretise age and hours-per-week to 5 categories for synthetic data generation, and converted back to continuous values for the downstream prediction, as our differentially private synthetic data generators require discrete data. We only remove rows with missing values in the included columns, so the number of rows is larger with the reduced set of features.
@@
(https://scikit-learn.org/stable/datasets/real_world.html#california-housing-dataset, license unknown) The california housing dataset contains information on housing districts, specifically census block groups, in California. The task is predicting the median house value in the district. We removed outlier rows where the average number of rooms is at least 50, or the average occupancy is at least 30. According to the dataset description, these likely correspond to districts with many empty houses. We log-transformed the target variable, as well as the population and median income features.
@@ E.2 Extra Plots
Figure S6: MSE prediction on the first two regression datasets. The predictions for synthpop are very accurate, and the predictions for DDPM are accurate for most cases. The linear regression measured MSE line for synthpop with ACS 2018 data is cut off due to excluding repeats with extremely large MSE ($\geq 10^{6}$). 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure S7: MSE prediction on the last two regression datasets. The predictions for synthpop are very accurate, and the predictions for DDPM are accurate for most cases. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
(c) German credit
Figure S8: Brier score prediction on three classification datasets. The predictions are accurate, but can have high variance. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.